OpenStudy (anonymous):

The sum of the digits of a two digit number is 12 the number formed by interchaning the digits is 54 more than the original number what is the original number? A 39 B48 C57

OpenStudy (bibby):

let's call the number ab where a is the tens digit and b is the ones digit

OpenStudy (bibby):

a+b=12 (the sum of the digits = 12)

OpenStudy (bibby):

and the number itself can be represented as a*10+b (a gets multiplied by 10 because it is the tens digit)

OpenStudy (bibby):

the interchanged number is b*10+a and we're told that the interchanged number is 54 more than the original so we have: a+b=12 10b+a=10a+b+54 that's a system of equations

OpenStudy (anonymous):

Wait where did the 54 come from? ._.

OpenStudy (bibby):

the question. do you know what interchanging is?

OpenStudy (anonymous):

I'm confused on exactly what it is. /).(\

OpenStudy (bibby):

no shame. basically if we have a number ab, the interchanged number is ba so if we have 36, the interchanged version would be 63

OpenStudy (bibby):

and so that's what I mean when I did the thing with times 10. If our number is 36, it can be written as \((3*10)+6\), aka \(36\) right? the interchanged number, 63, can be written as \((6*10)+3\) aka 63

OpenStudy (bibby):

I really suck at explaining this but this is the site I used to look this up http://www.onlinemathlearning.com/digit-problems.html

OpenStudy (anonymous):

I get a bit of what you're saying! :3 Thank you for helping me. :)

OpenStudy (bibby):

so can you solve the system?

OpenStudy (anonymous):

It would be 57 right?

OpenStudy (bibby):

well to check it, compare 57 and the interchanged value to start, what is 57 interchanged?

OpenStudy (anonymous):

5x10+7?

OpenStudy (bibby):

that's just 57, to interchange it, we flip the digits around

OpenStudy (anonymous):

OHHHH! So it would be 75?

OpenStudy (bibby):

yeahhh, and so now compare 57 and 75 is 75-57 = 54?

OpenStudy (anonymous):

noo.

OpenStudy (bibby):

which means 57 is not our answer if you're really lazy you could bruteforce your way through and just check each choice like that

OpenStudy (anonymous):

It'd be 39! :D

OpenStudy (bibby):

yeah, let me show you how it works if you solve the system tho

OpenStudy (anonymous):

okay :)

OpenStudy (bibby):

a+b=12 10b+a=10a+b+54 let's use substitution. we'll start by solving for a: \(a=12-b\) is that much clear?

OpenStudy (anonymous):

Mhmm.

OpenStudy (bibby):

we can then plug it into the second equation. \(10b+(12-b)=10(12-b)+b+54\) simplify that and we get \(9b+12=120-10b+b+54\\9b+12=174-9b\\18b=162\\b=\dfrac{162}{18}=9\)

OpenStudy (bibby):

and so that means that our "ones" digit has to be a 9. you can get the tens digit by solving for b and substituting

OpenStudy (anonymous):

Ohhh okay.

OpenStudy (bibby):

I can't seem to get a=3, but this calculator agrees with our answer http://www.wolframalpha.com/input/?i=a%2Bb%3D12+%2C10b%2Ba%3D10a%2Bb%2B54

OpenStudy (anonymous):

Okay! :D Thank you so much! <3

OpenStudy (bibby):

\(\huge \color{pink}{\heartsuit}\) <3